Operations Meeting August 3rd, 2009

  • 13:00 Eastern
  • Phone Number 510-665-5437 Meeting ID 786999
  • Attending: Rob, Tom, Burt, Scott Beardsley, Kyle, Doug, Elizabeth, Soichi, Levinth, Wayne Betts, Arvind, Karthik, Sam Hoover, Steve Timm

Operations Agenda

Operations Last Week

  • GOC sent a notification July 23rd requesting that OIM users login, and confirm their profile information -- please do this if you have not already!
  • BDII Review of Outage from July 8th Link
  • OSG 1.2.0: Released July 27th -
  • Previous Meeting Minutes

Operations This Week

Future Events

  • VORS Turn Down Scheduled for August 24th - This was extended while we contact the sites not reporting RSV to the GOC.
  • September Machine Room Move in Bloomington September 19 2009 - All GOC Services in Bloomington are expected to be down. IUPUI will still be handling BDII traffic. GOC will attempt to install as many other services on an Indianapolis based server/VM as possible, especially other critical ones like MyOSG?.

New Registrations

Name Support Center ResURL Site_Verify Status Other Info Ticket
AGLT2_CE_2 USATLAS   Passing RSV    

Removed from OSG

OSG RA Items & CA Questions

GOC Tickets

Relevant URLs

OSG Tickets

Ticket # Support Center Action/Summary Last Action Date GOCAssignee Notes

WLCG Tickets

7218 dq2-get failure
7148 440 errors at BU_ATLAS_Tier2 and increasing followup
5059 VOInfo entry should specify the SE it belongs to -- ISSUE=5059
7239 source error BNL-OSG2_MCDISK
7265 Open Science Grid: BDII Issue ISSUE=7145 PROJ=71
7019 CALTECH T2 FileOpenError? 'dcap:/
7246 Open Science Grid: BDII Issue ISSUE=7145 PROJ=71
7166 Some OSG sites claim to be in Antarctica


-- KyleGross - 03 Aug 2009

Topic revision: r4 - 03 Aug 2009 - 17:21:51 - ElizabethChism
Hello, TWikiGuest


TWiki | Report Bugs | Privacy Policy

This site is powered by the TWiki collaboration platformCopyright by the contributing authors. All material on this collaboration platform is the property of the contributing authors..